OpenStudy (anonymous):

8(x − 1) = 6(x + 2) − 4

OpenStudy (anonymous):

8(x − 1) = 6(x + 2) − 4 Distribute: 8x - 8 = 6x +12 - 4 then simplify 8x - 8 = 6x + 8 then + 8 to both sides ......... 8x = 6x + 16 the take 6X from both sides...2X = 16 so x = 8